OpenStudy (anonymous):

4 plus the fraction 2 which is over x+1

OpenStudy (anonymous):

What is the goal of the problem?

OpenStudy (anonymous):

addition of rational expressions :/

OpenStudy (anonymous):

okay, so I would multiply the 4 by (x+1)/(x+1) and the fraction by (4/4) then add the two together.

OpenStudy (anonymous):

ok so what i got then was 4x+12 over 4x+1

OpenStudy (anonymous):

remember your distributive property. is 4(x+1) = 4x + 1?

OpenStudy (anonymous):

oh shoot! so 4x+4 not 1

OpenStudy (anonymous):

exactly, also, can we reduce this new fraction any further?

OpenStudy (anonymous):

can't you eliminate the 4x's and you're left with 12/4 which will equal 3?

OpenStudy (anonymous):

we cannot actually do this, due to the rules of fractions. for example, say we had: (4 + 10) / (4 + 6) doing simple math we can simply this to: 14/10 and if we try to do what you are saying: ( 0 + 10 ) / ( 0 + 6 ) (crossing off the "4"s) we get 10/6 14/10 is not equal to 10/6 and thus this is not a valid math rule :/ If I were you, though, I would try to factor out any constants, if possible.

OpenStudy (anonymous):

Aw shoot :/ I am just not a math wiz at all :( so the lowest we can get the problem then would have to be just 4x+12/4x+4 ?

OpenStudy (anonymous):

I do see one common constant between the numerator and the denominator, perhaps we could factor that out and cancel?

OpenStudy (anonymous):

hmmm 4x+12 / 4x+4 = x+3